OpenStudy (anonymous):

determine the abs extreme values of the function f(X)=sinx - cosx+6 on the interval 0 to 2pi

OpenStudy (anonymous):

i got the derivative but dont know how to solve for x.

OpenStudy (anonymous):

You have to use extreme value theorem.

OpenStudy (anonymous):

because you're looking for absolute extremes, you'll need to look at the local max/min and the values at the endpoints. then you must compare those values which is the most extreme.

OpenStudy (anonymous):

what's your derivative?

OpenStudy (anonymous):

f'(x)=cosx+sinx

OpenStudy (anonymous):

ok, set f'(x) = 0.... and solve for x. is this where you're stuck?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i get -tanx=0

OpenStudy (anonymous):

\[\large cosx+sinx=0\rightarrow sinx=-cosx\rightarrow \frac{sinx}{cosx}=-1\rightarrow tanx=-1 \]

OpenStudy (anonymous):

where is tangent negative? what quadrants?

OpenStudy (anonymous):

kk got it. 2nd and 4th

OpenStudy (anonymous):

ok... what's x in those quadrants?

OpenStudy (anonymous):

i got it now.

OpenStudy (anonymous):

thanks.

OpenStudy (anonymous):

ok... we done then?

OpenStudy (anonymous):

yw... :)

OpenStudy (anonymous):

1 more.

OpenStudy (anonymous):

y=cos^3x

OpenStudy (anonymous):

wht u need done here?

OpenStudy (anonymous):

y'

OpenStudy (anonymous):

and equation of tanget at 3pi.

OpenStudy (anonymous):

chain rule... \[\large y=cos^3x=(cosx)^3 \] \[\large y'=3(cosx)^2(-sinx)=-3sinxcos^2x \]

OpenStudy (anonymous):

3pi = pi so \[\large y'|_{x=pi}=-3sin(\pi)cos^2(\pi)=-3(0)(-1)^2=0 \] this is the slope of the tangent line at x=pi

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

now i know the slope.. which is 0, how do i find the b value? (y=mx+b)